OpenStudy (loser66):

Let C be the circle \(x^2+y^2 =1\) oriented counterclockwise in the xy-plane. What is the value of the line integral \(\oint_C(2x-y)dx +(x+3y)dy \)

OpenStudy (loser66):

@dan815

ganeshie8 (ganeshie8):

you don't wanto use green's thm ?

OpenStudy (loser66):

If it helps, why not?

ganeshie8 (ganeshie8):

use it then

OpenStudy (loser66):

ok, let me try. Actually, I didn't know what the notation \(\oint\) mean. I never see it before. :)

ganeshie8 (ganeshie8):

that just means the curve is a closed loop

OpenStudy (loser66):

I am working on it, will tag you to check it later, ok?

ganeshie8 (ganeshie8):

ok, just need to find the curl and setup double integral

ganeshie8 (ganeshie8):

knw how to find the curl ?

OpenStudy (loser66):

yes, I divide it into 2 parts, \(y = \pm \sqrt{1-x^2}\) ,hence the limit for the first part will go from 0 to 1, right?

OpenStudy (loser66):

oh, We talk about 2 different things. ha!!

ganeshie8 (ganeshie8):

lol yeah actually we don't need to do much work here, find the curl, you will know why :)

OpenStudy (loser66):



OpenStudy (loser66):

ok, give me your way, please. hehehe..

ganeshie8 (ganeshie8):

find the curl first

ganeshie8 (ganeshie8):

\(\large Mdx + Ndy\) curl = \(N_x - M_y\)

ganeshie8 (ganeshie8):

\(\large (2x-y)dx +(x+3y)dy\) curl = ?

OpenStudy (loser66):

It looks like differential equation part? finding exactness, right?

ganeshie8 (ganeshie8):

\(\large (2x-y)dx +(x+3y)dy\) \(M = 2x-y\) \(N = x+3y\) \(N_x = 1\) \(M_y = -1\) curl = \(N_x - M_y = 1-(-1) = 2\)

OpenStudy (loser66):

I DO lost. :)

ganeshie8 (ganeshie8):

Easy.. just take partials and subtract

OpenStudy (loser66):

I know, but don't know why we have to do that.

ganeshie8 (ganeshie8):

because we want to use green's thm

OpenStudy (loser66):

I was taught that I have to find parametric equations for x, y and replace and take a loooooooooong steps to get the answer. This is somehow different.

OpenStudy (loser66):

But that is the reason i post the problem here to learn the shorter way. :)

ganeshie8 (ganeshie8):

\[\oint_C(2x-y)dx +(x+3y)dy ~~=~~ \iint_R~2 dxdy = 2\iint_R~1 dxdy = ?\]

OpenStudy (loser66):

You still use x, y , not r and theta?

OpenStudy (loser66):

oh, that is perimeter of the circle?

ganeshie8 (ganeshie8):

I just applied green's theorem to convert line integral into double integral

OpenStudy (loser66):

Ok, I got you. Thanks a lot. Need practice more. :)

OpenStudy (loser66):

One more question:

OpenStudy (loser66):

If the curve is not a circle, we must define the limits of x,y to put into the double integral, right?

ganeshie8 (ganeshie8):

green's theorem works only if the curve is a closed loop

OpenStudy (loser66):

Yes, Again, don't we have to change to polar form?

ganeshie8 (ganeshie8):

for all other cases you need to work it by parameterizing

OpenStudy (loser66):

YES.

ganeshie8 (ganeshie8):

you can but its not really needed here if you recall the fact that \(\iint_R ~1 ~dxdy\) represents the area of the region.

ganeshie8 (ganeshie8):

\[\oint_C(2x-y)dx +(x+3y)dy ~~=~~ \iint_R~2 dxdy = 2\color{red}{\iint_R~1 dxdy }= ?\] that red part represents the area of the circular unit disk

OpenStudy (loser66):

hey, on the previous comment (and you delete it), you stated the result is 4pi, ha!! now it turns to 2pi??

ganeshie8 (ganeshie8):

that red part is 2pi final answer is 4pi

OpenStudy (loser66):

how?

OpenStudy (loser66):

area of unit circle is pi

ganeshie8 (ganeshie8):

Oops! you're right haha

OpenStudy (loser66):

hihihi... ok, got you now. Much appreciate for being patient to me.

ganeshie8 (ganeshie8):

np :) maybe for practice, work it by parameterizing also

OpenStudy (loser66):

Yes, Sir.